Given a stock solution of yellow #5 with a concentration of 2.9 x 10-4 M, calculate the concentration of yellow #5 in the 4 prepared solutions (A thru D) assuming you perform these dilutions exactly as required in the procedure of this lab.

50 mL Beaker Dye stock solution 1:1 Ethanol/Water solvent
A 1.00 mL 9.00 mL
B 3.00 mL 7.00 mL
C 5.00 mL 5.00 mL
D 7.00 mL 3.00 mL
Answer #1

The common way to make a solution is to dilute a stock solution with additional solvent. The relationship between the concentrations and volumes of the concentrated and diluted solutions is given below.

CiVi = CfVf

where Ci and Vi are the concentration and volume of the initial (concentrated) solution and Cf and Vf are the concentration and volume of the final (diluted) solution.

A. 2.9 x 10-4 M * 1.00mL = Cf*10.00mL

Cf = 2.9 x 10-4 M * 1.00mL/10.00mL = 2.9 x 10-5 M

B. 2.9 x 10-4 M * 3.00mL = Cf*10.00mL

Cf = 2.9 x 10-4 M * 3.00mL/10.00mL = 8.7*10^-5 M

C. 2.9 x 10-4 M * 5.00mL = Cf*10.00mL

Cf = 2.9 x 10-4 M * 5.00mL/10.00mL =1.45 x 10-4 M

D. 2.9 x 10-4 M * 7.00mL = Cf*10.00mL

Cf = 2.9 x 10-4 M * 7.00mL/10.00mL = 2.03*10^-4 M